OpenStudy (anonymous):

The degree measure of one angle of a parallelogram is 30 more than that of a consecutive angle. Find the degree measure of each angle of the parallelogram. Info: The sum of the degree measures of any two consecutive angles of a parallelogram is 180*. ****Please show work so that I can understand.****

OpenStudy (anonymous):

consecutive angles are also called adjacent angles.......... In a parallelogram, consecutive/adjacent angles always add up to 180 degree (as they are co-interior angles of two parallel lines) Let the smaller angle be = x degree then larger angle is (x+30) degree so x + (x+30) = 180 2x + 30 = 180 2x = 180-30 x = 150/2 x = 75 so smaller angle is 75 degree and larger angle is 75+30 = 105 degree

OpenStudy (anonymous):

oh...thanks. It make sense that consecutive angles are ajacent. Thanks alot!